A gas thermometer measures temperature by measuring the pressure of a gas inside the fixed volume container a thermometer reads a pressure of 248 kPa at 0 degrees Celsius what is the temperature with thermometer reads a pressure of 345 kPa
1
Expert's answer
2020-03-05T07:57:21-0500

The thermometer is based on the isochoric process, which could be described with the following expression:


"\frac{P_{0}}{T_{0}}=\frac{P}{T};"


or "T=\frac{T_{0}P}{P_{0}};"


Using the given data:


"T=\frac{273K*345kPa}{248kPa}=380K(107C)."